OpenStudy (anonymous):

Complete the ordered pair so that it is a solution of 2/3 x + y - 1 = 0. (-6, ___) 3 4 5

OpenStudy (anonymous):

First rearrange the equation to solve for y given x, as follows: (2/3)(x) - 1 = -y 1 - (2/3)(x) = y then plug in x and solve: 1 - (2/3)(-6) = 5

OpenStudy (anonymous):

thank you so much